# A Comprehensive Survey of Contamination Detection Methods in Large Language Models

### v4

@@ 1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), large language models (LLMs) have emerged as a pivotal tool, driving innovation across a wide spectrum of applications, from natural language processing (NLP) and automated content creation (Achiam et al., 2023; Betker et al., 2023) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023b; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored writing (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination can significantly impair the models’ effectiveness, reliability, and generalization. In this study, we define the phenomenon of contamination as any leakage of the evaluation data within the training data.
@@
Detecting contamination in a LLM is a broad issue that we divide into open-data contamination detection and closed-data contamination detection depending on the nature of the training set. On the one hand, open-data refers to the scenario where the LLM pre-training data is known, enabling direct comparisons with the evaluation dataset. On the other hand, closed-data contamination refers to the more and more prevalent use case of an unknown pre-training set. This latter case is more challenging and necessitates an examination of the LLM’s behavior on evaluation data points. Contamination may span only part of a (x, y) data point where x is the input and y the corresponding label. If we consider whether the contaminated data contains only x or both x and y, contamination can be divided into input contamination and input+label contamination. Detecting contamination is further complicated by varying level of access to the LLM. With white-box access, the LLM and all weights are available, e.g., a local model. In gray-box access, model weights are not available to the researcher but output token probability distributions are. Most closed-source setups, e.g., GPT-4, follow a black-box setup where only the output text can be accessed from the API. We will review in this study contamination detection methods specifically tailored for each of these levels of model access.
@@
MIA, fill=purple!20, domain_box [SPV-MIA (Fu et al., 2023), CAMIA (Chang et al., 2024), S2-MIA (Li et al., 2024), EM-MIA (Kim et al., 2024), models ] ] ] [Model
@@
Memorization, fill=purple!20, domain_box [Elangovan et al. (2021); Magar & Schwartz (2022); Gemma Team Google DeepMind (2024), models ] ] [Prompt
@@
  • To the best of our knowledge, we are the first to thoroughly review the field of contamination detection in LLMs. We review all reported contamination detection methods up until early 2025, totaling more than 50 detection techniques and more than 100 relevant papers.
@@
We refer to Figure 1 for an overview of our proposed classification of contamination detection methods. The concurrent work of Palavalli et al. (2024) proposes a taxonomy of the different types of contamination. We review the detection of contamination at large, unrestricted to a specific form of contamination such as a particular noising then injection of test data points into the training set. The concurrent works of Deng et al. (2024) and Fu et al. (2024b) also survey contamination in LLMs. Compared to concurrent surveys, we dive deep with a unique focus on contamination detection methodologies and provide the most comprehensive overview of current papers, enabling enabling researchers to quickly grasp the existing research field of contamination detection in LLMs as of early 2025.
The rest of this paper is structured as follows: Section 2 explores methods and findings related to open-data contamination detection. Section 3 addresses the issue of closed-data contamination detection, where models may be evaluated on data that they have previously seen. Section 4 discusses the current and future challenges in the field, including best practices, and reviews all new datasets proposed to detect contamination.
@@ 2 Open-Data Contamination Detection
In the open-data setup, we assume that the (pre)-training set is known. Given a training dataset $D$ and an evaluation dataset $D_{E}$, the objective is to check whether $D\cap D_{E}$ is empty. In the following, we review existing open-data contamination detection methods, which directly compare ensembles of data points.
@@ 3 Closed-Data Contamination Detection
In closed-data contamination detection, the training data is unknown. Formally, a model $M$ has been trained on a training set $D_{M}$ (the pre-training set is specific to each LLM) and is applied to an evaluation dataset $D_{E}$. $D_{M}$ is not accessible, and researchers have to use tools analyzing $M$’s behavior on $D_{E}$.
@@ 3.2.3 Limitations of Membership-Inference Attacks
Membership-Inference attacks are hard in practice with neural networks. Duan et al. (2024) evaluate a series of MIAs on the Pythia LLMs (from 70M to 12B parameters), pre-trained on the Pile. Across types of attacks and domains, performance hardly surpasses random guessing (ROC-AUC inferior to 0.6). This failure is attributed to the common practice in LLM development of pre-training only for a single epoch and on a too large dataset, and the fact that the border between members and non-members is not clear. By splitting both groups based on time, where non-members are drawn from datasets posterior to the pre-training cutoff date, MIA accuracy drastically improves. This finding motivates the monitoring of LLMs performance through time described above. Meeus et al. (2024b) echo these findings from Duan et al. (2024): they train a trivial bag-of-words model on MIA datasets constructed by splitting on time and reach a high ROC-AUC. Authors argue that proper MIA evaluation should be done on a randomized train-test split. Concurrent work from Maini et al. (2025) also question the validity of membership inference attacks on a time-split of Wikipedia. An experiment replacing member sentences by sentences from the validation set of The Pile (Gao et al., 2020) using Pythia models (Biderman et al., 2023) shows membership inference tests still reaching an ROC-AUC of 0.7 ; proving that they merely capture a temporal distribution shift rather than actual membership to the pre-training set. Despite this lack of success of MIA on pre-training sets of LLMS, Li et al. (2024) show that similarity between the target sample and the generated output of a LLM-powered RAG system yields a powerful MIA on the RAG database. On Natural Questions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), the method reaches more than 0.85 in both ROC-AUC and PR-AUC.
@@ 3.3 Model Memorization
An intuitive approach to detect contamination in LLMs is to frame it as a memorization problem. While MIAs assess the presence of an entire data point into the training set, here memorization is defined in a larger sense and focuses on detecting memorized sequences of tokens which may not be full data points (documents), but rather subsets of them.
@@ 3.3.1 Quantifying Memorization
Elangovan et al. (2021) highlight that some LLMs may benefit from an overlap between training and test sets, and merely memorize data points in this intersection, inflating their test set performance. This overlap is due to the common practice of simple random data shuffling in NLP tasks. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection. Magar & Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. They propose to quantify the effect of contamination by training on a mix of general and task-specific data, then comparing performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation.
@@ 3.4.2 Probing Hidden Layers
Model Memorization Gemma Team Google DeepMind (2024)
Magar & Schwartz (2022)
@@
DICE (Tu et al., 2024)
Table 2: Classification of closed-data contamination detection methods as per the level of access to the LLM required.
Table 2 summarizes closed-data detection techniques organized by the minimum level of access to the LLM which is required. All methods based on performance analysis only require a black-box access, making them convenient to use. Membership-inference attacks are equally split between black-box and gray-box access requirements. Model memorization techniques mainly rely on prompting and thus only need black-box access. Lastly, for model confidence methods, which often analyze output distributions, a gray-box or even white-box model access is needed.
@@ 4.1 Best Practices to Avoid Contamination
Scanning newly released evaluation datasets
Sainz et al. (2023) argue to develop automatic or semi-automatic approaches to detect contamination for new benchmarks and design mechanisms to flag those works with contamination.
Evaluating on a wide spectrum
@@
Several data contamination mitigation techniques are detailed in the concurrent survey work of Xu et al. (2024a).
@@ 4.2 New Evaluation Benchmarks
Open-Source Pre-training Data
When the pre-training data is publicly available, the task of defining members and non-members of the pre-training set becomes drastically simpler. Duan et al. (2024) introduce the MIMIR dataset, constructed from The Pile training set (Gao et al., 2020), on the which popular open-source models such the Pythia (Biderman et al., 2023) and GPT-Neo [^2] https://github.com/EleutherAI/gpt-neo series are pre-trained. MIMIR covers domain-specific subsets (e.g. GitHub, Wikipedia, Arxiv) and extracts members through several n-gram sizes, allowing researchers to work on multiple levels of membership granularity. MIMIR has become widely used in MIA research with LLMs. OLMoMIA (Kim et al., 2024) introduces a membership inference dataset centered around the open-source OLMo-7B LLM (Groeneveld et al., 2024). Knowing that a pre-training epoch roughly consists of 450k, training data seen before 100k pre-training steps is considered as member, and training data 400k and 450k steps is non-member. Dolma-Book (Zhang & Wu, 2024) also leverages the publicly available pre-training data from OLMo, and samples non-member books from books from the Project Gutenberg [^3] https://www.gutenberg.org/ dated after January 1st, 2024.
@@
Some new datasets discard the static paradigm of existing LLM evaluation datasets, and instead propose a dynamic, periodically updated benchmark to avoid contamination:
@@
Alternative Evaluation
New approaches to LLM evaluation are emerging in the era of widespread public benchmark contamination. The DCQ benchmark utilizes GPT-4 to paraphrase data points from famous benchmarks such as WNLI (Levesque et al., 2012), and checks whether the LLM picks the exact continuation or a paraphrased version, where the former strongly suggests contamination. Moreover, Alzahrani et al. (2024) point out that slight perturbations to existing benchmarks may disrupt model rankings on leaderboards, calling for better evaluation benchmarks.

### v5

@@ 1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), large language models (LLMs) have emerged as a pivotal tool, driving innovation across a wide spectrum of applications, from natural language processing (NLP) and automated content creation (Achiam et al., 2023; Betker et al., 2023) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023b; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored writing (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination can significantly impair the models’ effectiveness, reliability, and generalization. Contamination refers to leakage of evaluation data within the training data, an undesirable phenomenon which will inflate the model’s performance on the evaluation data.
We identify several primary sources of such contamination that can arise during large-scale data collection and curation. First, unfiltered web scraping to collect pre-training data often introduces data with minimal oversight, increasing the likelihood of inadvertently capturing benchmark or test datasets that are publicly available online. Second, contamination may result from the unintentional inclusion of evaluation materials during preprocessing or aggregation of training corpora, particularly when datasets are not rigorously audited. Finally, the reuse of proprietary or copyrighted material without robust provenance tracking can also lead to inadvertent overlap between training and evaluation sets. These factors underscore the need for meticulous dataset design and transparent documentation practices to ensure the validity of experimental results.
@@
1.1 Contamination and Contamination Detection Types
Contamination may span only part of a (x, y) data point where x is the input and y the corresponding label. If we consider whether the contaminated data contains only x or both x and y, contamination can be divided into input contamination and input+label contamination.
Detecting contamination in a LLM is a broad issue that we divide into open-data contamination detection and closed-data contamination detection depending on the nature of the training set. On the one hand, open-data refers to the scenario where the LLM pre-training data is known, enabling direct comparisons with the evaluation dataset. On the other hand, closed-data contamination refers to the more and more prevalent use case of an unknown pre-training set. This latter case is more challenging and necessitates an examination of the LLM’s behavior on evaluation data points.
Detecting contamination is further complicated by varying level of access to the LLM. With white-box access, the LLM and all weights are available, e.g., a local model. In gray-box access, model weights are not available to the researcher but output token probability distributions are. Most closed-source setups, e.g., GPT-4, follow a black-box setup where only the output text can be accessed from the API. We will review in this study contamination detection methods specifically tailored for each of these levels of model access.
@@ 1.1 Contamination and Contamination Detection Types
MIA, fill=purple!20, domain_box [SPV-MIA (Fu et al., 2023), CAMIA (Chang et al., 2024), S2-MIA (Li et al., 2024b), EM-MIA (Kim et al., 2024), models ] ] ] [Model
@@
Memorization, fill=purple!20, domain_box [ Dynamic Soft Prompting (Wang et al., 2024),Elangovan et al. (2021); Magar & Schwartz (2022); Carlini et al. (2023); Gemma Team Google DeepMind (2024), models ] ] [Prompt
@@ 1.2 Contributions and Structure of this Survey
  • To our knowledge, we are one of the first to thoroughly review the field of contamination detection in LLMs, unrestricted to a specific type of methods or a specific performance aspect. We review all reported contamination detection methods up until early 2025, totaling more than 50 detection techniques and more than 120 relevant papers. We followed an inclusive approached and kept all open-access papers relevant to the problem of contamination detection in LLMs, regardless of publication status. The vast majority of resulting papers which are included in this survey are from 2023-2025.
@@
We refer to Figure 1 for an overview of our proposed classification of contamination detection methods.
The rest of this paper is structured as follows. Section 2 formally defines contamination and associated concepts. Section 3 explores methods and findings related to open-data contamination detection. Section 4 addresses the issue of closed-data contamination detection. Section 5 discusses the current and future challenges in the field, including best practices, and reviews all new datasets proposed to detect contamination.
1.3 Comparison to Related Surveys
There already exists surveys on the broad issue of contamination.
Deng et al. (2024) survey contamination in language models, covering the problem from detection, to its impact and methods of remediating it. Compared to their work, our survey solely focuses on detecting contamination, and covers a more exhaustive set of contamination detection techniques.
The work of Palavalli et al. (2024) proposes a taxonomy of the different types of contamination. In Figure 1, our second-level breakdown of closed-data contamination detection into dataset-level methods and instance-level methods echos their categories. Unlike their survey, we review contamination detection only, and we do so at large, unrestricted to a specific form of contamination such as a particular noising then injection of test data points into the training set.
Unlike the two surveys above, Fu et al. (2024b) focus on contamination detection, like our work. They also propose a taxonomy of detection methods, which has a few similarities with ours (like instance similarity, which we cover within open-data methods), but our classification is more fine-grained (four levels of classification, whereas their classification only has one). The authors review assumptions behind certain detection methods and challenge their true efficiency. Notably, they show that membership-inference attacks (MIAs) perform poorly in detecting pre-training contamination. Our study does not specifically focus on the success of contamination detection methods: we are rather interested in describing in depth the whole landscape of contamination detection, and explain how each method works, including requirements in terms of model access (Table 2) and best suited training stage (Table 3).
2 Formal Definitions
Formally, let $M$ be a model (e.g. Llama-3 (Dubey et al., 2024)) and $\mathcal{D}$ be a data distribution on an input space $\mathcal{X}$. We denote by $D_{M}$ the dataset that $M$ was pre-trained on (e.g, The Pile (Gao et al., 2020)), and $D_{\text{eval}}$ an evaluation dataset of interest (e.g, MMLU (Hendrycks et al., 2020)).
Definition 1 (Contamination)
We say that the dataset $D_{\text{eval}}$ is contaminated by the dataset $D_{M}$ if:
$\exists x\in D_{\text{eval}},\exists x^{\prime}\in D_{M},f(x,x^{\prime})>=\tau$ (1)
where $f$ is a similarity function, and $\tau$ a similarity threshold. $\tau=1$ corresponds to verbatim or strict contamination, but we do not restrict ourselves to this setup in this paper.
Contamination signs often include inflated performance on evaluation benchmarks, and model memorization of entire substrings. Contamination detection refers to techniques assessing whether Equation 1 holds true or not.
Definition 2 (Membership Inference Attack)
A membership inference attack (MIA) is a function attempting to assess whether a data point was part (a member of) the training set $D_{M}$, and can be noted as:
$A:\mathcal{X}\times M(\mathcal{X})\rightarrow\{0,1\}$ (2)
where
$A(x,M(x))$ $=1\quad\text{if }x\in D_{M}$
$A(x,M(x))$ $=0\quad\text{otherwise}$
Both concepts of contamination detection and membership inference are closely related yet fundamentally different: a membership inference attack can be used to detect contamination, but membership inference attacks can also be defined and succeed without the existence of contamination.
Definition 3 (Model Memorization)
Let us note $G(p)$ the model’s natural language output on natural language prompt $p$. We say that $M$ has memorized a data point $x\in D_{M}$ if:
$\Pr[G(p_{x})=x\mid x\in D_{M}]\gg\Pr[G(p_{x})=x\mid x\sim\mathcal{D}]$ (3)
for some prompt $p_{x}$ that elicits $x$ (typically, $p_{x}$ is a truncation of $x$).
In other words, memorization refers to the LLM’s capability to recreate a training data verbatim when prompted accordingly.
@@ 3 Open-Data Contamination Detection
In the open-data setup, we have access to the (model-specific) pre-training set $D_{M}$. In this case, contamination detection falls back to investigating the intersection $D_{M}\cap D_{\text{eval}}$ for any evaluation dataset $D_{\text{eval}}$. In the following, we review existing open-data contamination detection methods, which directly compare ensembles of data points.
@@ 4 Closed-Data Contamination Detection
In closed-data contamination detection, the pre-training set $D_{M}$ is unknown. Since $D_{M}$ is not accessible, researchers have to use tools analyzing the model’s behavior on the evaluation dataset $D_{\text{eval}}$.
@@ 4.2.3 Limitations of Membership-Inference Attacks
Membership-Inference attacks are hard in practice with neural networks. Duan et al. (2024) evaluate a series of MIAs on the Pythia LLMs (from 70M to 12B parameters), pre-trained on the Pile. Across types of attacks and domains, performance hardly surpasses random guessing (ROC-AUC inferior to 0.6). This failure is attributed to the common practice in LLM development of pre-training only for a single epoch and on a too large dataset, and the fact that the border between members and non-members is not clear. By splitting both groups based on time, where non-members are drawn from datasets posterior to the pre-training cutoff date, MIA accuracy drastically improves. This finding motivates the monitoring of LLMs performance through time described above. Meeus et al. (2024b) echo these findings from Duan et al. (2024): they train a trivial bag-of-words model on MIA datasets constructed by splitting on time and reach a high ROC-AUC. Authors argue that proper MIA evaluation should be done on a randomized train-test split. Concurrent work from Maini et al. (2025) also question the validity of membership inference attacks on a time-split of Wikipedia. An experiment replacing member sentences by sentences from the validation set of The Pile (Gao et al., 2020) using Pythia models (Biderman et al., 2023) shows membership inference tests still reaching an ROC-AUC of 0.7 ; proving that they merely capture a temporal distribution shift rather than actual membership to the pre-training set. Despite this lack of success of MIA on pre-training sets of LLMS, Li et al. (2024b) show that similarity between the target sample and the generated output of a LLM-powered RAG system yields a powerful MIA on the RAG database. On Natural Questions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), the method reaches more than 0.85 in both ROC-AUC and PR-AUC.
@@ 4.3 Model Memorization
An intuitive approach to detect contamination in LLMs is to frame it as a memorization detection problem. While MIAs assess the presence of an entire data point into the training set, here memorization is more granular and detects memorized sequences of tokens which may not be full data points (documents), but rather subsets of them. The rate of memorized data points can be used to assess the presence of contamination. For this endeavor, it is important to first be able to quantify memorization itself.
@@ 4.3.1 Quantifying Memorization
Elangovan et al. (2021) highlight that some LLMs may benefit from an overlap between training and test sets, and merely memorize data points in this intersection, inflating their test set performance. This overlap is due to the common practice of simple random data shuffling in NLP tasks. By employing a bag-of-words approach to assess text similarity, the study provides a foundational method for identifying and mitigating leakage, though it acknowledges the complexity of semantic similarity measurement and the potential for more sophisticated methodologies to refine overlap detection. Magar & Schwartz (2022) find evidence of contamination across several benchmarks, impacting the performance of models like GPT-3. They propose to quantify the effect of contamination by training on a mix of general and task-specific data, then comparing performance on seen versus unseen instances to measure memorization and exploitation. This methodology highlights the nuanced relationship between memorization and exploitation. Carlini et al. (2023) propose a simple method to detect memorization by measuring the completion rates of randomly sampled prefixes from the training set, and discuss the relations between memorization with model capacity, data repetition, context size. Memorization is more widespread than previously thought, and model size, data frequency and prompt length are factors leading to an increase in memorization.
@@
As we have seen in Definition 3, detecting memorization sums to finding a (prefix, suffix) mapping such that the corresponding suffix can be perfectly completed when conditioning the model on the prefix. For example, Özdayi et al. (2023) append a learnable soft prompt penalized through gradient ascent to prevent memorizing the original suffix. Wang et al. (2024) employs a transformer-based generator to approximate such mapping of prefixes, and use it to estimate the memorization rate on unseen corpus.
@@ 4.4.2 Probing Hidden Layers
Model Memorization Magar & Schwartz (2022)
Carlini et al. (2023)
Gemma Team Google DeepMind (2024)
Dynamic Soft Prompting Wang et al. (2024)
@@
Oren et al. (2023)
DICE (Tu et al., 2024)
Liu et al. (2024)
Table 2: Classification of closed-data contamination detection methods as per the level of access to the LLM required. Black-box methods simply need to prompt the model, while gray-box methods require access to the output tokens distributions and white-box techniques need full access to model weights.
4.5 Practical Usage of Closed-Data Contamination Detection Methods
In the previous subsections, we reviewed the different types contamination detection methods. We now wish to give more concrete intuition in how these methods can be applied in practice.
Contamination detection techniques strongly differ in terms of model access requirement.Table 2 summarizes closed-data detection techniques organized by the minimum level of access to the LLM which is required. We define three levels of model access: black-box, which merely needs prompting access ; gray-box, which needs access to output tokens distributions ; and white-box which needs to see model weights. All methods based on performance analysis only require a black-box access, making them convenient to use. Membership-inference attacks are equally split between black-box and gray-box access requirements. Model memorization techniques mainly rely on prompting and thus only need black-box access. Lastly, for model confidence methods, which often analyze output distributions, a gray-box or even white-box model access is needed.
Closed-Data Method Type Contamination Detection Method Pre-training Instruction-tuning RLHF
Performance Analysis CAP Zhao et al. (2024)
Aiyappa et al. (2023)
Roberts et al. (2023)
Li & Flanigan (2023)
Cao et al. (2024)
Membership Inference Attacks RECALL (Xie et al., 2024)
MaxNorm(Meeus et al., 2024a)
PREMIA (Feng et al., 2024)
MIA-Tuner (Fu et al., 2024a)
SPV-MIA (Fu et al., 2023)
CAMIA (Chang et al., 2024)
EM-MIA (Kim et al., 2024)
S2-MIA (Li et al., 2024b)
Model Memorization Magar & Schwartz (2022)
Carlini et al. (2023)
Gemma Team Google DeepMind (2024)
Dynamic Soft Prompting Wang et al. (2024)
Nasr et al. (2023)
According To prompting (Weller et al., 2023)
DCQ (Samuel et al., 2024)
Local order quiz (Samuel et al., 2024)
Chang et al. (2023)
TS-Guessing (Deng et al., 2023)
Ranaldi et al. (2024)
Model Confidence Min-KProb (Shi et al., 2024)
DC-PDD (Zhang et al., 2024b)
Min-K%++ (Zhang et al., 2024a)
@@ 4.5 Practical Usage of Closed-Data Contamination Detection Methods
Table 3: Classification of closed-data contamination detection techniques according to the model training stage(s) which they are more suitable to be applied to.
We also highlight that closed-data contamination detection techniques may naturally be applied to different model training stages. In Table 3, we indicate the most suitable training stage(s) where each method can be applied, among (i) pre-training, (ii) instruction-tuning and (iii) reinforcement learning from human feedback (RLHF). Prompt tuning is naturally tailored for RLHF models which can be prompted with any user query. Cloze tasks and likelihood-based methods are suitable for pre-trained models, which have captured tokens distributions. We emphasize that this categorization is flexible and not strictly defined: for instance, methods most suited for pre-trained models, like Min-KProb, can in theory be applied to instruction-tuned or RLHF models. However, results should be cautiously interpreted as RLHF often smoothens overconfident predictions to align with human preferences, weakening the contamination signal.
@@ 5.1 Best Practices to Avoid Contamination
Scanning newly released evaluation datasets
Sainz et al. (2023) argue to develop automatic or semi-automatic approaches to detect contamination for new benchmarks and design mechanisms to flag those works with contamination. This process goes hand in hand with creating new, non-contaminated evaluation datasets.
Evaluating on a wide spectrum
@@
Several data contamination mitigation techniques are detailed in the survey work of Xu et al. (2024a).
5.2 Towards Open-Source Pre-Training Data
As we have seen in this study, access to the pre-training data is the first fundamental characterization of contamination detection technique. When the pre-training data are publicly available, the task of defining membership to the pre-training set becomes drastically simpler. To facilitate contamination detection, it is important for the research community to move towards the general usage of open-source pre-training sets.
Duan et al. (2024) introduce the MIMIR dataset, constructed from The Pile training set (Gao et al., 2020), on the which popular open-source models such as the Pythia (Biderman et al., 2023) and GPT-Neo [^2] https://github.com/EleutherAI/gpt-neo model series are pre-trained. MIMIR covers domain-specific subsets (e.g. GitHub, Wikipedia, Arxiv) and extracts members through several n-gram sizes, allowing researchers to work on multiple levels of membership granularity. MIMIR has become widely used in MIA research with LLMs. OLMoMIA (Kim et al., 2024) introduces a membership inference dataset centered around the open-source OLMo-7B LLM (Groeneveld et al., 2024). Knowing that a pre-training epoch roughly consists of 450k, training data seen before 100k pre-training steps is considered as member, and training data 400k and 450k steps is non-member. Dolma-Book (Zhang & Wu, 2024) also leverages the publicly available pre-training data from OLMo, and samples non-member books from books from the Project Gutenberg [^3] https://www.gutenberg.org/ dated after January 1st, 2024.
@@ 5.3 New Evaluation Benchmarks
Benchmark Perturbation
Several new datasets are crafted by perturbing existing benchmarks, for instance through rephrasing instructions or adding more complex instructions. Perturbed data points are more challenging for LLMs, and can be designed to be adversarial examples aiming to fool the LLMs. Moreover, Alzahrani et al. (2024) point out that slight perturbations to existing benchmarks may disrupt model rankings on leaderboards, calling for better evaluation benchmarks. Such benchmarks include:
  • GSM-Plus (Li et al., 2024a). Using the 1,319 questions of the grade-school level maths word problems GSM8K dataset (Cobbe et al., 2021) as seed dataset, the authors apply eight perturbations of five different types: numerical variation, arithmetic variation, problem understanding, distractor insertion and critical thinking. The dataset is generated with GPT-4 (manually verified by humans) and has 10,552 questions in total. High relative performance drop rates of above 30% are observed for open-source models. Several elaborate prompting techniques such as chain-of-thought fall short of providing meaningful improvement, but a custom compositional prompting technique provides notable performance improvement.
  • MATH-Perturb (Huang et al., 2023) introduces two flavors of variations to 279 problems from the hardest level (level 5) of MATH dataset (Hendrycks et al., 2021): slightly different problems solvable with the same method (MATH-P-Simple), and problems needing a different approach (MATH-P-HARD). Performance degrades notably for all LLMS on MATH-P-HARD. Interestingly, a lot of errors can be traced to memorization: the model simply blindly applies the same technique used to solve the original problem.
@@
Some new datasets discard the static paradigm of existing LLM evaluation datasets, and instead propose a dynamic, ever-changing or periodically updated benchmark to avoid contamination.
  • LiveBench (White et al., 2024) is a challenging benchmark made of 18 tasks from 6 categories: math, coding, reasoning, language comprehension, instruction following, and data analysis. Tasks are more complicated versions of existing benchmarks, such as BigBench-Hard (Srivastava et al., 2022), and are evaluated against an objective ground-truth rather than with human or LLM preference. To avoid contamination, one sixth of questions are replaced every month, chosen among oldest and easiest tasks.
  • Chatbot Arena (Zheng et al., 2023) is a popular crowdsourced LLM evaluation platform. Humans submit queries and are shown answers from two LLMs which quality they have to rank. By construction, this benchmark is not set in stone and changes LLM rankings as more users interact with the platform. Besides, the design of Chatbot Arena limits contamination as user-generated prompts and model response rankings are less likely to be trained on.
Several dynamic benchmarks are specifically crafted for membership-inference attacks: